OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4).

OpenStudy (anonymous):

y = four thirdsx + 8 y = four thirdsx + 3 y = -three fourthsx + 8 y = -three fourthsx + 3

OpenStudy (anonymous):

4x-3y=-12 4x+12=3y y=4/3x+4 what i have so far: 

OpenStudy (anonymous):

we need the same slope, just a different y-intercept. so if the point given is 4 down, we also do -4 to the y-intercept. so the new equation is